# Methods Paper Version 19

### v1

@@ IV.2 Two-channel results
Figure 13: a) ${}^{3}S_{1}$ phase shifts $\delta_{0}$, b) ${}^{3}D_{1}$ phase shifts $\delta_{2}$, c) ${}^{3}SD_{1}$ mixing parameter $\epsilon$ obtained by the HORSE method with $\hbar\Omega=30$ MeV and various $N_{\textrm{max}}$; the smoothing parameter $a=2.5$. The inset in (c) displays $\epsilon$ in a larger scale on the y-axis.
We present in Fig. 13 the results of calculations of the phase shifts and mixing parameter $\epsilon$ in the ${{}^{3}SD}_{1}$ coupled partial waves of $np$ scattering using the HORSE method with $\hbar\Omega=30$ MeV and smoothing parameter $a=2.5$. The ${}^{3}S_{1}$ phase shifts $\delta_{0}$ convergence is fast and at $N_{\max}=6$ we have accurate results. The ${}^{3}D_{1}$ phase shift $\delta_{2}$ convergence is also good and starting from $N_{\max}=6$ the phase shifts are reasonable (note the expanded scale in Fig. 13b as compared with the scale in Fig. 13a) while with ${N_{\max}=10}$ and $N_{\max}=12$, we obtain an excellent description of the phase shifts. The mixing parameter $\epsilon$ is also reasonably well-described at $N_{\max}=12$ (note, the scale in Fig. 13c is further expanded), we obtain small deviations of $\epsilon$ from the exact values at $N_{\max}=16$ and a very accurate description of the mixing parameter at $N_{\max}=20$.
@@
If we use a “complete” set of SRFs $\{\beta_{q}\}$ in Eq. (1) which is enough to form a complete basis in the subspace of the oscillator functions $\phi_{ni}(\vec{r})$ with $n=0,1,...\,,{\cal N}_{i}$, $i=1,2$, the Efros approach is equivalent to the HORSE method. This set of SRFs $\{\beta_{q}\}$ may be arbitrary: it is only important that the functions $\beta_{q}$ are linearly-independent. In this case, one can use the over-complete Eq. (27) which suggest additional relations for independent calculations of the unknowns that should be equal such as $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)$ in the case of two-channel scattering. As an example we consider the $S$-matrix calculations at the energy of $E=0.5$ MeV in the coupled ${}^{3}SD_{1}$ partial waves of $np$ scattering with ${\hbar\Omega=30}$ MeV and smoothing parameter $a=2.5$. We employ the set of $H^{tr}$ eigenfunction SRFs for $N_{\textrm{max}}=12$ and $N=15$, i. e., all of the 13 $H^{tr}$ eigenfunctions are used in the set of the SRFs $\{\beta_{q}\}$ and two additional oscillator functions used exclusively for orthogonalization are included in the set $\{\bar{\beta}_{q}\}$ as in Eq. (77). We obtain the $S$ matrix
@@
As it was already noted, in the above case of the “complete” set of SRFs $\{\beta_{q}\}$, the Efros and HORSE methods are equivalent. Since the “complete” set of SRFs in current ab initio nuclear structure applications is extremely large, it is more interesting to investigate the essential simplification suggested within the Efros method as compared to the HORSE method which makes it possible to use a reduced set of SRFs. Let us consider the same case as above, but for $N=12$. In this case, Eq. (77) results in the $S$ matrix
@@
Figure 14: a) ${}^{3}S_{1}$ phase shifts $\delta_{0}$, b) ${}^{3}D_{1}$ phase shifts $\delta_{2}$, c) ${}^{3}SD_{1}$ mixing parameter $\epsilon$ obtained by the Efros method with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and with smoothing parameter ${a=2.5}$ using the eigenfunction SRFs and ${}^{3}S_{1}$ as the first channel with $i=1$, i. e., imposing the symmetry by setting $\mathbb{S}_{12}(E)$ equal to $\mathbb{S}_{21}(E)$. See text for a detailed explanation of the label for each curve.
Figure 15: The same as Fig. 14 but using the ${}^{3}D_{1}$ as the first channel with $i=1$.
In the case of a limited set of SRFs, if we use the ${}^{3}S_{1}$ as the first channel corresponding to $i=1$ and set $l_{1}=0$ and $l_{2}=2$ [i. e., first solving the complete set of the ${}^{3}S_{1}$ channel equations, obtaining $\mathbb{S}_{21}(E)$, and next solving the reduced set of the ${}^{3}D_{1}$ equations imposing the symmetry by setting $\mathbb{S}_{12}(E)$ equal to $\mathbb{S}_{21}(E)$], we get results different from those obtained using the ${}^{3}D_{1}$ as the first channel corresponding to $i=1$ and setting $l_{1}=2$ and $l_{2}=0$ (i. e., solving first the complete set of equations in the ${}^{3}D_{1}$ partial wave and next switching to the reduced set of equations in the ${}^{3}S_{1}$ partial wave). We compare these approaches in Figs. 14 and 15 where we present, respectively, the results using the ${}^{3}S_{1}$ and the ${}^{3}D_{1}$ channel as the first.
The results for the phase shifts and the mixing parameter $\epsilon$ presented in Figs. 14 and 15 were obtained using $H^{tr}$ eigenfunctions as SRFs, $\hbar\Omega=30$ MeV, and truncation of the Hamiltonian $H^{tr}$ at $N_{\max}=12$ with the smoothing parameter $a=2.5$. We prefer to use the lowest-lying eigenfunctions of $H^{tr}$ in calculations. Let us introduce the notation used in the label for Fig. 14 (which is the same as for Fig. 15) by way of an explicit example. The $N=5\,\{0,1,2\}$ in the figure legends means that we use 5 unknowns in the $i=1$ partial wave [3 SRF coefficients $b^{2}_{q}(E)$, $q=0,1,2$ plus 2 $S$-matrix elements $\mathbb{S}_{11}(E)$ and $\mathbb{S}_{21}(E)$], 4 unknowns in the $i=2$ partial wave [3 SRF coefficients $b^{1}_{q}(E)$, $q=0,1,2$ plus 1 $S$-matrix element $\mathbb{S}_{22}(E)$], that is a total of $\color[rgb]{0,0,0}\mathbb{N}=9$ independent unknowns, and take as the SRFs the lowest (ground state) eigenfunction $\beta_{0}^{\textrm{Efn}}(\vec{r})$, $\beta_{1}^{\textrm{Efn}}(\vec{r})$, and $\beta_{2}^{\textrm{Efn}}(\vec{r})$. We omit ${\beta}_{2}^{\textrm{Efn}}(\vec{r})$ in the set $\{\bar{\beta}_{q}^{2}\}$ used for the orthonormalization of the reduced set of equations in the $i=2$ channel, but not in the set $\{\bar{\beta}_{q}^{1}\}$ used in the equations in the $i=1$ channel.
In the case of using the ${{}^{3}S}_{1}$ as the first channel presented in Fig. 14, the calculations with $N=5$ provide reasonable results for $\color[rgb]{0,0,0}\delta_{0}$ at low energies and for $\color[rgb]{0,0,0}\epsilon$ in a larger energy interval but a completely incorrect behavior of $\color[rgb]{0,0,0}\delta_{2}$. It is interesting to note that the results for $\delta_{0}$ at $N=5$ start deviating from the exact curve at energies around $E=11{-}12$ MeV. This roughly corresponds to the second excited state eigenenergy of $H^{tr}$ at $E_{2}=11.457$ MeV. The origin of strange unreasonable $\delta_{2}$ behavior is unclear. In the case of $N=6$, $\delta_{0}$ describes well the exact result while $\color[rgb]{0,0,0}\delta_{2}$ and $\color[rgb]{0,0,0}\epsilon$ are reasonably accurate up to approximately $E=25$ MeV and $E=20$ MeV, respectively. In calculations with ${N=7}$ and $8$, we see emerging numerical problems. For example, a jump of $\delta_{0}$ around ${E=3}$ MeV occurs which is more pronounced in the $N=7$ than in the $N=8$ calculations. Except for these spikes, the $\delta_{0}$ convergence is reasonable. The calculations of $\delta_{2}$ with $N=7$ fail completely and with $N=8$ fail up to approximately 12 MeV but describe well the exact results at larger energies. The $\epsilon$ calculations with $N=7$ and $8$ provide close results which are reasonable at energies above 16 MeV but unreasonable at lower energies.
More reasonable results are obtained if setting the ${}^{3}D_{1}$ as the first channel (see Fig. 15). With the same parameters ($\hbar\Omega$, $N_{\max}$, etc.) we get good results with $N=5$ and $N=6$ at least up to $E=16$ MeV for all quantities. The results are still accurate in cases of $N=7$ and $N=8$, though $\color[rgb]{0,0,0}\delta_{0}$ obtained with $N=8$ shows some oscillation around $E\approx 12$ MeV and $\color[rgb]{0,0,0}\delta_{2}$ and $\color[rgb]{0,0,0}\epsilon$ have jumps around $\color[rgb]{0,0,0}E\approx 3$ MeV in cases of both $N=7$ and $N=8$.
Figure 16: ${}^{3}SD_{1}$ channel: a) $|\textrm{det}(A)|$, b) $\textrm{Arg}(\textrm{det}(A))$ obtained in calculations with $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and the ${{}^{3}D}_{1}$ partial wave as the first channel with and without the basis rearrangement. The vertical line corresponds to the first excited state eigenenergy of $\color[rgb]{0,0,0}H^{tr}\!$.
Figure 17: a) ${}^{3}S_{1}$ phase shifts $\delta_{0}$, b) ${}^{3}D_{1}$ phase shifts $\delta_{2}$, c) ${}^{3}SD_{1}$ mixing parameter $\epsilon$ obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and with smoothing parameter ${a=2.5}$ using the eigenfunction SRFs and ${}^{3}S_{1}$ as the first channel with $i=1$ and rearranging the set of SRFs to eliminate the vanishing point in the determinant of $A$. This elimination is checked over a range of $E$ from 0 to 50 MeV.
Figure 18: The same as Fig. 17 but using ${}^{3}D_{1}$ as the first channel with $i=1$.
Figure 19: The same as Fig. 18 but calculations are performed with $N_{\max}=20$.
Figure 20: Absolute magnitudes of ${{}^{3}SD_{1}}$ wave function amplitudes at the energy of $E=0.5$ MeV obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and ${}^{3}D_{1}$ as the first channel with $i=1$: a) $|d_{n,22}(E)|$ (“strong” ${}^{3}S_{1}$ component), b) $|d_{n,21}(E)|$ (“weak” ${}^{3}S_{1}$ component), c) $|d_{n,12}(E)|$ (“weak” ${}^{3}D_{1}$ component), d) $|d_{n,11}(E)|$ (“strong” ${}^{3}D_{1}$ component).
Figure 21: Arguments of ${{}^{3}SD_{1}}$ wave function amplitudes at the energy of $E=0.5$ MeV obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and ${}^{3}D_{1}$ as the first channel with $i=1$: a) $\textrm{Arg}(d_{n,22}(E))$ (“strong” ${}^{3}S_{1}$ component), b) $\textrm{Arg}(d_{n,21}(E))$ (“weak” ${}^{3}S_{1}$ component), c) $\textrm{Arg}(d_{n,12}(E))$ (“weak” ${}^{3}D_{1}$ component), d) $\textrm{Arg}(d_{n,11}(E))$ (“strong” ${}^{3}D_{1}$ component).
Figure 22: $\frac{3}{2}^{-}$ resonance in $n\alpha$ scattering supported by the Woods–Saxon potential: a) resonance energy $E_{r}$ and b) resonance width $\Gamma$ obtained by locating the $S$-matrix pole within the HORSE method using various $N_{\max}$ truncations, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$.
Figure 23: $\frac{3}{2}^{-}$ resonance in $n\alpha$ scattering supported by the Woods–Saxon potential: a) resonance energy $E_{r}$ and b) resonance width $\Gamma$ obtained by locating the $S$-matrix pole within the Efros method with eigenfunction SRFs using various $N$, $N_{\max}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$ in comparison with the exact values and with the HORSE method results.
Figure 24: Deuteron bound state energy $E_{b}$ obtained with Daejeon16 $NN$ interaction using various $N_{\textrm{max}}$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$ by the HORSE method and by the direct diagonalization of $H^{tr}$.
Figure 25: Deuteron bound state energy $E_{b}$ obtained with Daejeon16 $NN$ interaction by locating the $S$-matrix pole via the Efros method using the eigenfunction SRFs, various $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$ in comparison with the exact value and with the HORSE method result.
Figure 26: Absolute magnitudes of expansion coefficients $|\breve{d}_{n,i}(E_{b})|$ of the deuteron wave function obtained by locating the $S$-matrix pole within the Efros method with different $N$ using eigenfunction SRFs, $N_{\max}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$: a) ${}^{3}S_{1}$ components, b) ${}^{3}D_{1}$ components. The deuteron energy obtained with each $N$ is given in the legend in parentheses. We present also for comparison the exact values of $|\breve{d}_{n,i}(E_{b})|$ ($E_{b}=-2.225$ MeV) and those values obtained by the HORSE method. Symbols are often obscured by other symbols especially in panel (a).
Figure 27: a) Energy $E_{r}$ and b) width $\Gamma$ of the resonance in the two-channel Noro–Taylor model Norro1980 in HORSE calculations with $\Omega=1.5$ au, smoothing parameter $a=5$, and various $N_{\max}$.
Figure 28: a) Energy $E_{r}$ and b) width $\Gamma$ of the resonance in the two-channel Noro–Taylor model Norro1980 in calculations within Efros method with eigenfunction SRFs, $\Omega=1.5$ au, $N_{\max}=20$, smoothing parameter $a=5$, and various $N$ as compared with the exact results and with the HORSE method results.
@@
The origin of these spikes is illustrated in Fig. 16, where $|\det(A)|$, the absolute magnitude of the matrix $A$ determinant, and $\textrm{Arg}\big{(}\textrm{det}(A)\big{)}$, the determinant argument, are plotted as functions of $E$ in the case of $N=7$ and using the ${}^{3}D_{1}$ as the first channel. If we use the set of the SRF eigenfunctions with increasing eigenvalues, $N=7\,\{0,1,2,3,4\}$, there is a clearly seen discontinuity of $\textrm{Arg}\big{(}\det(A)\big{)}$ around $E=3$ MeV which coincides with the spikes in Fig. 15 for $\color[rgb]{0,0,0}\delta_{2}$ and $\color[rgb]{0,0,0}\epsilon$. To fix this, we simply replace one of the SRFs $\beta_{q}^{\textrm{Efn}}(\vec{r})$ and change their ordering which results in replacing the omitted function $\bar{\beta}_{q}^{\textrm{Efn}}(\vec{r})$ in the orthonormalization of the $i=2$ equations. For example, the ordering $\{0,1,4,6,2\}$ fixes the discontinuities in $\textrm{Arg}(\textrm{det}(A))$ and the spikes that they cause.
This adjustment of the set of SRFs improves the results as demonstrated in Figs. 17 and 18 where modified sets of SRFs are indicated by the red colors in the legends. It is seen that with $N=8\,\{0,1,2,6,4,3\}$ in the case of ${{}^{3}S}_{1}$ used as the first channel or with $N=8\,\{0,1,2,3,4,6\}$ in the case of ${{}^{3}D}_{1}$ used as the first channel, we obtain an excellent description of both phase shifts $\delta_{0}$ and $\delta_{2}$ but an inaccurate description of the mixing parameter $\epsilon$. It is interesting that in the case of the first channel assigned to the ${{}^{3}S}_{1}$ partial wave, the best $\epsilon$ description is obtained with the smallest number of SRFs ${N=5\,\{0,2,1\}}$ when the phase shifts $\delta_{0}$ and $\delta_{2}$ are well-described in relatively small energy intervals, up to approximately 15 and 30 MeV, respectively (see Fig. 17). It is seen that the convergence trends with an increasing number of SRFs $N$ is complicated and non-monotonic in the two-channel case. We have not yet found a systematic way to change the SRF set ordering to eliminate the vanishing of the $A$ determinant. Generally, our preference is to retain the lowest-lying eigenfunctions since they are important in the formation of scattering observables at low energies.
As was already mentioned, the Efros method with a “complete” set of SRFs becomes equivalent to the HORSE approach. The description of the phase shifts with increasing $N$ approaches, though non-monotonically, to that of the HORSE method. As is seen in Fig. 13, the HORSE description of the scattering parameters, in particular, of the mixing parameter $\epsilon$, is still not ideal at $N_{\max}=12$ with $\hbar\Omega=30$ MeV. Therefore, to obtain a good description of all scattering observables of $np$ scattering in the ${{}^{3}SD}_{1}$ coupled partial waves in a wide energy range, one needs to perform calculations with a larger $N_{\max}$ value. We present in Fig. 19 the results of calculations with $N_{\max}=20$, with the rest of the parameters the same as in Fig. 18. It is seen that we get an excellent description of all scattering observables up to the energy of 50 MeV if we use a large enough set of SRFs $N=11\,\{0,1,2,3,4,5,6,7,8\}$.

### v2

@@ I Naive application
Figure S3: $^{1}S_{0}$ phase shifts obtained by the HORSE method for various $N_{\textrm{max}}$ with $\hbar\Omega=30$ MeV.
Figure S4: $^{1}S_{0}$ phase shifts obtained by the HORSE method for various $\hbar\Omega$ at $N_{\textrm{max}}=24$.
Figure S5: The effect of smoothing with various parameters $a$ on calculations of the $^{1}S_{0}$ phase shifts by the HORSE method with $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV.
Figure S6: $^{1}S_{0}$ phase shifts obtained by the HORSE method with various $N_{\textrm{max}}$ at $\hbar\Omega=30$ MeV with smoothing with parameter $a=2.5$.
However in the case of an “incomplete” set of harmonic oscillator SRFs $\bar{\beta}^{\rm HO}_{q}(\vec{r})$ with $N<\mathcal{N}$, by solving Eq. (33), we obtain a set of functions $g_{n}(E)$ with $n=0$, … , N2N-2 in Eq. (52) and get nonzero expansion coefficients $d_{n}(E)=a^{as}_{n}(E)$} with $n\geq N-1$ which mimic the role of the last three terms in Eq. (56) though these terms were not explicitly included in $H^{tr}$.
II Single-channel scattering
Figure S7: $^{1}S_{0}$ phase shifts obtained by the adapted Efros method using a) $N=4$ and $N_{\textrm{max}}=10$, b) $N=4$ and $N_{\textrm{max}}=14$}, and c) $N=5$ and $N_{\textrm{max}}=14$} with both the harmonic oscillator and eigenfunction SRFs, $\hbar\Omega=30$} MeV.
Figure S8: $^{1}S_{0}$ phase shifts obtained by the adapted Efros method with various NN at $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV using the eigenfunction SRFs.
Figure S9: $^{1}S_{0}$ phase shifts obtained by the adapted Efros method using $N=8$, $N_{\textrm{max}}=14$, and $\hbar\Omega=30$} MeV using the eigenfunction SRFs without smoothing and with smoothing at $a=2.5$.
Figure S10: Coefficients $d_{n}(E)$ of the wave function expansion (51) at energy E=0.5{E=0.5} MeV in the $^{1}S_{0}$ partial wave obtained with various NN at $N_{\textrm{max}}=14$} and $\hbar\Omega=30$ MeV with smoothing parameter $a=2.5$ using the eigenfunction SRFs.
Figure S11: $^{1}S_{0}$ phase shifts obtained by the adapted Efros method with various $q_{0}$ using the hybrid SRF set given by Eq. (S8), $N_{\textrm{max}}=14$, $N=6$, $\hbar\Omega=30$ MeV, and the smoothing parameter $a=2.5$. Vertical lines show the energies of eigenstates of the Hamiltonian $H^{tr}$.
Figure S12: The same as Fig. S12 but using the hybrid SRF set given by Eq. (S9) and $N_{\max}=12$.
In this section, we start by discussing the convergence of single-channel scattering calculations using the HORSE method. Figure S3 shows the convergence of the $^{1}S_{0}$} npnp scattering phase shifts supported by the Daejeon16 $NN$ interaction with $N_{\textrm{max}}$ at $\hbar\Omega=30$ MeV. It is seen that at $N_{\textrm{max}}=18$, we obtain a very good description of the phase shifts while at $N_{\textrm{max}}=24$, a complete convergence is achieved in a wide range of energies.
Figure S4 shows the change of the phase shift description with $\hbar\Omega$ at $N_{\textrm{max}}=24$. We see that $\hbar\Omega=30$ MeV is optimal from the point of view of convergence of calculations.
Figure S5 demonstrates the effect of smoothing on the phase shift calculations within the HORSE method with $N_{\textrm{max}}=14$. We see that our choice of using the parameter values of $a=2.5$ or $a=5$ in our studies of $NN$ scattering is justified.
In Fig. S6, we find that excellent convergence of the phase shifts within the HORSE method is achieved with smoothing at $N_{\textrm{max}}=14$ as opposed to the $N_{\textrm{max}}=24$ needed for convergence without smoothing.
[…]
Figure S8 shows the convergence of $\delta_{0}(E)$ obtained by the adapted Efros method with $N$ for $N_{\textrm{max}}=14$. We note a reasonable convergence at $N=5$. However, at larger $N$, namely at $N=8$, the results start deviating significantly from the exact result. The irregularity of convergence is a consequence of the short-range behavior of the $NN$ interaction and a sudden drop-off of its matrix elements $V_{nn^{\prime}}$ at $n,n^{\prime}=\cal N$ due to the sharp truncation at relatively small $N_{\textrm{max}}$.
Figure S13: a) $^{3}S_{1}$ phase shifts $\delta_{0}$, b) $^{3}D_{1}$ phase shifts $\delta_{2}$, c) $^{3}SD_{1}$ mixing parameter $\epsilon$ obtained by the HORSE method with $\hbar\Omega=30$ MeV and various $N_{\textrm{max}}$; the smoothing parameter $a=2.5$. The inset in (c) displays $\epsilon$ in a larger scale on the y-axis.
Figure S14: a) $^{3}S_{1}$ phase shifts $\delta_{0}$, b) $^{3}D_{1}$ phase shifts $\delta_{2}$, c) $^{3}SD_{1}$ mixing parameter $\epsilon$ obtained by the adapted Efros method with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and with smoothing parameter a=2.5{a=2.5} using the eigenfunction SRFs and $^{3}S_{1}$ as the first channel with $i=1$, i. e., imposing the symmetry by setting $\mathbb{S}_{12}(E)$ equal to $\mathbb{S}_{21}(E)$.
Figure S9 demonstrates the effect of smoothing with parameter $a=2.5$ on the phase shift calculations within the adapted Efros method with $N_{\textrm{max}}=14$, $N=8$. We see complete convergence when smoothing is used in contrast to when it is not used.
Figure S10 shows the convergence with $N$ of the wave function expansion coefficients $d_{n}(E)$ [see Eq. (51)] at the energy E=0.5{E=0.5} MeV where the $^{1}S_{0}$} phase shift obtained with smoothing at $N_{\max}=14$ and $\hbar\Omega=30$ MeV is close to the exact. The oscillator amplitudes $d_{n}(E)$ are seen to converge reasonably well.
We consider also a couple of sets of nonorthonormal linearly independent SRFs combining both the harmonic oscillator and the eigenfunctions,
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...\,,q_{0},\\ \phi_{\mathcal{N}-q+q_{0}}(\vec{r}),&q=q_{0}+1,...\,,N-2,\\ \phi_{\mathcal{N}}(\vec{r}),&q=N-1,\end{cases}$ (S8)
and
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...\,,q_{0},\\ \phi_{q-q_{0}-1}(\vec{r}),&q=q_{0}+1,...\,,N-2,\\ \phi_{\mathcal{N}}(\vec{r}),&q=N-1.\end{cases}$ (S9)
Here, $q_{0}+1$ is the number of low-lying eigenfunctions used. For both Eqs. (S8) and (S9), the set of SRFs $\{\beta_{q^{\prime}}\},$ $q^{\prime}=0,...\,,N-2$ appearing explicitly in Eq. (19) is the same as the set of SRFs $\{\bar{\beta}_{q}\}$ for q=0,,N2q=0,...\,,N-2 but it excludes $\bar{\beta}_{N-1}=\phi_{\mathcal{N}}(\vec{r})$.
In Fig. S12 we present the $^{1}S_{0}$ phase shifts obtained with the SRF set of Eq. (S8) with $N_{\textrm{max}}=14$, $\hbar\Omega=30$ MeV, and $N=6$. Vertical lines correspond to the energies of eigenstates of the Hamiltonian $H^{tr}$ at which the phase shifts are well-described. We see that at least $q_{0}=3$ is needed to get reasonable phase shifts in a wide range of energies while with $q_{0}=4$ the results for the phase shifts are excellent.
Figure S12 presents the same results but obtained with the SRF set of Eq. (S9) and smaller $N_{\textrm{max}}=12$}. We see that in this case, we get reasonable results even with $q_{0}=0$ and much more accurate with $q_{0}=1$. Thus it is preferable to adopt the SRF set of Eq. (S9).
Figure S15: $^{3}SD_{1}$ channel: a) $|\textrm{det}(A)|$, b) $\textrm{Arg}(\textrm{det}(A))$ obtained in calculations with $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and the ${{}^{3}D}_{1}$ partial wave as the first channel with and without the basis rearrangement. The vertical line corresponds to the first excited state eigenenergy of $\color[rgb]{0,0,0}H^{tr}\!$.
Figure S16: a) $^{3}S_{1}$ phase shifts $\delta_{0}$, b) $^{3}D_{1}$ phase shifts $\delta_{2}$, c) $^{3}SD_{1}$ mixing parameter $\epsilon$ obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and with smoothing parameter a=2.5{a=2.5} using the eigenfunction SRFs and $^{3}S_{1}$ as the first channel with $i=1$ and rearranging the set of SRFs to eliminate the vanishing point in the determinant of $A$. This elimination is checked over a range of $E$ from 0 to 50 MeV.
III Two-channel scattering
We present in Fig. S13 convergence with $N_{\textrm{max}}$ of the phase shifts and mixing parameter $\epsilon$ in the ${{}^{3}SD}_{1}$ coupled partial waves of $np$ scattering using the HORSE method and smoothing with the parameter $a=2.5$. We see that $N_{\textrm{max}}=12$ gives reasonable results for both phase shifts but still some small deviations from the exact results for the mixing parameter $\epsilon$ while the complete convergence is in a wide energy range is achieved at $N_{\max}=20$.
@@ III Two-channel scattering
If we use a “complete” set of SRFs $\{\beta_{q}\}$ in Eq. (1) which is enough to form a complete basis in the subspace of the oscillator functions $\phi_{ni}(\vec{r})$ with $n=0,1,...\,,{\cal N}_{i}$, $i=1,2$, the adapted Efros method is equivalent to the HORSE method. This set of SRFs $\{\beta_{q}\}$ may be arbitrary: it is only important that the functions $\beta_{q}$ are linearly independent. In this case, one can use the over-complete Eq. (27) which suggest additional relations for independent calculations of the unknowns that should be equal such as $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)$ in the case of two-channel scattering. As an example we consider the $S$-matrix calculations at the energy of $E=0.5$ MeV in the coupled $^{3}SD_{1}$ partial waves of $np$ scattering. We employ the set of $H^{tr}$ eigenfunction SRFs for $N_{\textrm{max}}=12$ and $N=15$, i. e., all of the 13 $H^{tr}$ eigenfunctions are used in the set of the SRFs $\{\beta_{q}\}$ and two additional oscillator functions used exclusively for orthogonalization are included in the set $\{\bar{\beta}_{q}\}$ as in Eq. (68). We obtain the $S$ matrix
@@
As it was already noted, in the above case of the “complete” set of SRFs $\{\beta_{q}\}$, the adapted Efros and HORSE methods give the same results. Since the “complete” set of SRFs in current ab initio nuclear structure applications is extremely large, it is more interesting to investigate the essential simplification suggested within the adapted Efros method as compared to the HORSE method which makes it possible to use a reduced set of SRFs. Let us consider the same case as above, but for $N=12$. In this case, Eq. (68) results in the $S$ matrix
@@
Figure S17: a) $^{3}S_{1}$ phase shifts $\delta_{0}$, b) $^{3}D_{1}$ phase shifts $\delta_{2}$, c) $^{3}SD_{1}$ mixing parameter $\epsilon$ obtained by the adapted Efros method with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and $^{3}D_{1}$ as the first channel with $i=1$, i. e., imposing the symmetry by setting $\mathbb{S}_{21}(E)$ equal to $\mathbb{S}_{12}(E)$.
In the case of the adapted Efros method with a limited set of SRFs, we present in Fig. S14 the phase shift convergence in which we use the $^{3}S_{1}$ as the first channel corresponding to $i=1$ and set $l_{1}=0$ and $l_{2}=2$. We use the same notation in Fig. S14 as that used in the label for Fig. 2 in the main text. In general, we see reasonable results for $N=5$ and $N=6$ at certain energy intervals but emerging numerical problems for N=7{N=7} and $8$ in the form of jumps.
Figure S18: The same as Fig. 2 in the main text, but calculations are performed with $N_{\max}=20$.
@@
As illustrated in Fig. S15, the origin of these spikes is due to the vanishing determinant of $A$ given by Eq. (35) at the energy $E$ around 3 MeV. Here, we plot $|\det(A)|$, the absolute magnitude of the matrix $A$ determinant, and $\textrm{Arg}\big{(}\textrm{det}(A)\big{)}$, the determinant argument, are plotted as functions of $E$ in the case of $N=7$ and using the $^{3}D_{1}$ as the first channel. If we use the selection of SRF eigenfunctions with increasing eigenvalues, $N=7\,\{0,1,2,3,4\}$, there is a clearly seen discontinuity of $\textrm{Arg}\big{(}\det(A)\big{)}$ around $E=3$ MeV which coincides with the spikes in Fig. S14 for $\color[rgb]{0,0,0}\delta_{2}$ and $\color[rgb]{0,0,0}\epsilon$. The SRF selection $\{0,1,4,6,2\}$ fixes the discontinuities in $\textrm{Arg}(\textrm{det}(A))$ and the spikes that they cause. In particular, we simply replace one of the SRFs $\beta_{q}^{\textrm{Efn}}(\vec{r})$ and change their selection which results in replacing the omitted function $\bar{\beta}_{q}^{\textrm{Efn}}(\vec{r})$ in the orthonormalization of the $i=2$ equations. We have not yet found a systematic way to change the SRF selection to eliminate the vanishing of the $A$ determinant. Generally, our preference is to retain the lowest-lying eigenfunctions since they are important in the formation of scattering observables at low energies. This adjustment of the set of SRFs improves the results as demonstrated in Fig. S16. It is seen that the convergence trends with an increasing number of SRFs $N$ can be complicated and non-monotonic in the two-channel case.
Figure S19: Absolute magnitudes of $^{3}SD_{1}$} wave function amplitudes at the energy of $E=0.5$ MeV obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and $^{3}D_{1}$ as the first channel with $i=1$: a) $|d_{n,22}(E)|$ (“strong” $^{3}S_{1}$ component), b) $|d_{n,21}(E)|$ (“weak” $^{3}S_{1}$ component), c) $|d_{n,12}(E)|$ (“weak” $^{3}D_{1}$ component), d) $|d_{n,11}(E)|$ (“strong” $^{3}D_{1}$ component).
Figure S20: Arguments of $^{3}SD_{1}$} wave function amplitudes at the energy of $E=0.5$ MeV obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and $^{3}D_{1}$ as the first channel with $i=1$: a) $\textrm{Arg}(d_{n,22}(E))$ (“strong” $^{3}S_{1}$ component), b) $\textrm{Arg}(d_{n,21}(E))$ (“weak” $^{3}S_{1}$ component), c) $\textrm{Arg}(d_{n,12}(E))$ (“weak” $^{3}D_{1}$ component), d) $\textrm{Arg}(d_{n,11}(E))$ (“strong” $^{3}D_{1}$ component).
In Fig. S17, we present the convergence of the phase shifts and mixing angle with $N$ at $N_{\textrm{max}}=12$ obtained using the $^{3}D_{1}$ as the first channel corresponding to $i=1$ and setting $l_{1}=2$ and $l_{2}=0$ (i. e., solving first the complete set of equations in the $^{3}D_{1}$ partial wave and next switching to the reduced set of equations in the $^{3}S_{1}$ partial wave). We obtain reasonable results for $N=5$ and $N=6$ at least up to $E=16$ MeV for all quantities. The results are still accurate in cases of $N=7$ and $N=8$. However, $\delta_{0}$ obtained with $N=8$ shows some oscillation around $E\approx 12$ MeV and $\delta_{2}$ and $\color[rgb]{0,0,0}\epsilon$ have jumps around $E\approx 3$ MeV in cases of both $N=7$ and $N=8$. Those numerical issues are fixed by adopting the SRF selection in Fig. 2 in the main text.
Figure S18 shows the results of calculations with $N_{\max}=20$}, with the rest of the parameters the same as in Fig. 2 in the main text. It is seen that we get an excellent description of all scattering observables up to the energy of 50 MeV if we use a large enough set of SRFs $N=11\,\{0,1,2,3,4,5,6,7,8\}$}.